A circular pond has a diameter of 2.5 meters. The owner wants to put a concrete walkway around the pool that has uniform width. She has enough to cover 5m(squared). What is the maximum width her walkway can be?
Answer by khwang(438)

If the width of the walkway is x meters, then its area is
pi(x + 2.5/2)^2 - pi (2.5/2)^2 = pi(x^2 + 2.5 x)
Solve: pi(x^2 + 2.5 x) <= 5,
or x^2 + 2.5 x - 5/pi <= 0
use pi = 3.1416 and quadratic formula,
two roots of x^2 + 2.5 x - 5/pi = 0 are
0.525963319, -3.025963319
when x is between these two roots, x^2 + 2.5 x - 5/pi <= 0
but x width shouldbe > 0.
Hence,x <= 0.5259 meters... max possible value of width
Kenny